# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ 1.3. Proof Scheme
Recall that Tverberg’s theorem [32] asserts that any set of $n\geq(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ can be partitioned into $r$ pairwise disjoint sets whose convex hulls have non-empty $r$-fold intersection. The $r=2$ case is Radon’s theorem. For prime powers, one has a topological extension of Tverberg theorem [22, 33] which asserts that any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ from the $(n-1)$-dimensional simplex to $\mathbb{R}^{d}$ has an $r$-Tverberg tuple – that is, an $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ of non-empty pairwise disjoint faces of $\Delta_{n-1}$ such that $f(\sigma_{1})\cap\cdots\cap f(\sigma_{r})\neq\emptyset$. When $r$ is a prime power, Tverberg’s theorem is recovered by letting $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be a linear map, that is, one which sends each convex combination of vertices of the simplex to the corresponding convex combination of the image of the vertices. In what follows, we identify the vertex set of $\Delta_{n-1}$ with $[n]=\{1,\ldots,n\}$ and we let $\operatorname{\mathrm{Vert}}(\sigma)\subseteq[n]$ denote the vertex set of any non-empty face $\sigma$ of $\Delta_{n-1}$.
@@
Let $r\geq 2$ be a prime power, and let $d,m,n\geq 1$ be integers with $n\geq(r-1)(d+m+1)+1$. Suppose that $\mathcal{F}$ is a family of subsets of $[n]$ such that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Then for any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$, there exists an $r$-Tverberg-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that no element of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{j})$ for any $j\in[r]$.
As a further example of fan distributions obtainable from Tverberg-type results, we show that the Gale dual of the of the “optimal colored Tverberg theorem” gives rainbow distributions of a coloring of a finite point set, that is an $r$-fan which distributes the point set so that each of the fan’s open half-flats contains at most one point from each color class.
@@
    Let $r\geq 4$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(d+1)$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{i}|\geq r$ for all $i\in[d+1]$, then there exists an $r$-fan of codimension $r-2$ which rainbow distributes the $(d+1)$-coloring.
@@
    Let $r\geq 2$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(2d+1)-1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{2d+1}$ is a $(2d+1)$-coloring of $X$ such that $|X_{i}|\geq r$ for all $i\in[2d+1]$, then there exists a complex regular $r$-fan in $\mathbb{C}^{n-d-1}$ which rainbow distributes the $(2d+1)$-coloring.
As an example of Theorem 1.6, again let $r=4$ and $d=1$. Then for any partition of a set of $8$ affinely spanning points in $\mathbb{R}^{6}$ into two color classes of size $4$ each, there is a $4$-dimensional $4$-fan in $\mathbb{R}^{6}$ such that the interior of each half-flat contains at most one point from each color class. Moreover, if $X$ is typical then the interiors of the fan’s half-flats must share at least $7$ of the $8$ points.
@@ 1.4. Distributions by Multiple Fans
Arguably the most fundamental problem in equipartition theory, dating back to Grünbaum [14] and Hadwiger [15], seeks the minimum dimension $d:=\Delta(m,k)$ such that any $m$ finite point sets in $\mathbb{R}^{d}$ can be equipartitioned by the $2^{k}$ regions determined by $k\geq 2$ independent hyperplanes. While a degrees of freedom argument  [25] gives the lower bound $\Delta(m,k)\geq\frac{m(2^{k}-1)}{k}$, very few tight upper bounds are currently known. For two hyperplanes in particular, sharp results exist only when either $m-1$, $m$, or $m+1$ is a power of two [7, 6, 19], in which case $\Delta(m,2)=\lfloor 3m/2\rfloor$. When $r$ is an odd prime, we give analogous extensions of our equidistribution results for two fans, where now given an $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a finite point set $X$ in $\mathbb{R}^{d}$ we ask for two $r$-fans $F_{r}^{1}=\cup_{i=1}^{r}B^{1}_{i}$ and $F_{r}^{2}=\cup_{j=1}^{r}B^{2}_{j}$ in $\mathbb{R}^{d}$ whose intersection contains $X$ so that each intersection of the form $\operatorname{\mathrm{Int}}(B_{i}^{1})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})$ contains at most $1/r^{2}$ of the points from each $X_{k}$. As with the calculations of the tight values of $\Delta(m,2)$, our results are ultimately derived from cohomological techniques which impose restrictions on the values of $m$.
@@
Let $d\geq 1$ be an integer, let $r\geq 3$ be a prime number, and let $m\geq 1$ be an integer such that each of the coefficients in the $r$-ary expansion of $m(r-1)/2$ is even.
@@
    If $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$, then any $m$-coloring of a set $X$ of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by two $r$-fans of codimension $r-2$.
@@
It is straightforward to show (see Remark 6.2) that the $r$-array expansion of $m(r-1)/2$ has all even coefficients provided $m$ is of the form $m=2a(r^{\ell_{1}}+\cdots+r^{\ell_{k}})$ where the $\ell_{i}\geq 0$ are distinct and $11\leq a\leq r-$ 1 is odd. In particular, one may let $m=2a(r^{\ell}-1)/(r-1)$ for any $\ell\geq 1$.
As was the case for a single fan, Theorem 1.7 is the special case of a Dolnikov-type piercing distribution result for two fans (see Theorem 8.1 below) which in turn is the Gale dual of a linear version of a multiple intersection version of Sarkaria’s theorem. In essence, this result guarantees two Tverberg pairs which are maximally distinct (see  [12] for the corresponding extension of Theorem 1.5 when $r=2$).
@@
Let $d,m,r,n\geq 1$ be integers where $r\geq 3$ is a prime number, the $r$-ary coefficients of $m(r-1)/2$ are all even, and $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$. Suppose that $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and that $\mathcal{F}$ is a family of non-empty subsets of $[n]$ such that $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$. Then there exists two Tverberg-tuples $(\sigma_{1},\ldots,\sigma_{r})$ and $(\tau_{1},\ldots,\tau_{r})$ for $f$ such that no member of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{i}\cap\tau_{j})$ for any $i,j\in[r]$.
@@ 2. Real and Complex Gale Duality
In this section we discuss the properties of the real and complex Gale transform which we will need in translating between Tverberg partitions and fan distributions. These are quite standard in the real setting (see, e.g., [20]) and only require the minor adjustment of taking conjugates to hold in the complex setting. Nonetheless, for the sake of completeness we shall include their proofs here.
@@
Now let $\mathbb{K}$ denote the field of real or complex numbers and let $a_{1},\ldots,a_{n}$ be a sequence of $n$ points in $\mathbb{K}^{d}$ which ($\mathbb{K}$-) affinely span $\mathbb{K}^{d}$ (thus $n\geq d+1$). The Gale transform of this sequence is obtained as follows. First, let $A$ be the $(d+1)\times n$ matrix with coefficients in $\mathbb{K}$ whose column vectors are the $\widehat{a_{j}}=(a_{j},1)\in\mathbb{K}^{d+1}$. Because the $a_{j}$ affinely span $\mathbb{K}^{d}$, the $\widehat{a_{j}}$ linearly span $\mathbb{K}^{d+1}$. Thus $A$ has maximal rank and so its kernel has dimension $n-d-1$. Letting $b_{1},\ldots,b_{n-d-1}$ be a basis for $\ker A$, we let $B$ be the $(n-d-1)\times n$ matrix whose $i$-th row is $b_{i}$, and we let $\overline{B}$ denote its conjugate matrix. The Gale transform of $a_{1},\ldots,a_{n}$ is then defined to be the columns $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ of $\overline{B}$. While the Gale transform is defined only up to a choice of basis for $\ker A$, the basic properties we shall need hold independently of such a choice, and moreover is not difficult to show that a different choice of basis corresponds to a change of the $g_{i}$ by a linear isomorphism of $\mathbb{K}^{n-d-1}$.
@@
Now suppose that $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ satisfies (1) and (2). Let $B$ be the $(n-d-1)\times n$ matrix whose columns are the $g_{i}$ and let $b_{1},\ldots,b_{n-d-1}$ be its rows. As the $g_{i}$ linearly span $\mathbb{K}^{n-d-1}$, $B$ has maximum rank and so $\dim\ker B=d+1$. Since $\sum_{i}g_{i}=0$, one can choose a basis $w_{1},\ldots,w_{d+1}$ for $\ker B$ with $w_{d+1}$ the all one vector. Now define $A$ to be the matrix whose rows are the $w_{i}$ and consider its conjugate $\bar{A}$. The columns of $A$ span $\mathbb{K}^{d+1}$ and therefore so do those of $\bar{A}$. As the last row of $\bar{A}$ is the all 1 vector, the columns of $\bar{A}$ are of the form $\widehat{a_{1}}=(a_{1},1),\ldots,\widehat{a_{n}}=(a_{n},1)$. Thus $a_{1},\ldots,a_{n}$ affinely span $\mathbb{K}^{d}$. By construction, we have $\ker\bar{A}=(\text{Row}\,A)^{\perp}=(\ker B)^{\perp}=\text{Row}\,\overline{B}.$ Thus $\overline{b}_{1},\ldots,\overline{b}_{n-d-1}$ is a basis for $\ker\bar{A}$, and so the columns $g_{1},\ldots,g_{n}$ of $\overline{\overline{B}}=B$ are the Gale transform of $a_{1},\ldots,a_{n}$. ∎
For our purposes, the essential property of the Gale transform is that it gives a correspondence between any affine dependency among the $a_{1},\ldots,a_{n}$ and an allocation of the $g_{1},\ldots,g_{n}$ with respect to a codimension one linear subspace. More precisely, let $\langle u,w\rangle=\sum_{i=1}^{n}u_{i}\overline{w}_{i}$ denote the standard Hermitian inner product of two vectors $u$ and $w$ in $\mathbb{K}^{n}$. One then has the following.
@@ 3. Tverberg Partitions and Linear Fans
In this section, we describe an explicit connection between Tverberg $r$-tuples and linear $r$-fans, i.e, those whose center is a linear affine space. This connection will be used crucially in the proofs of Theorems 1.11.31.4, and 1.6 in Section 4.
@@
3.1. Real Linear Fans

### v3

@@ 1.3. Proof Scheme
Recall that Tverberg’s theorem [35] asserts that any set of $n\geq(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ can be partitioned into $r$ pairwise disjoint sets whose convex hulls have non-empty $r$-fold intersection. The $r=2$ case is Radon’s theorem. For prime powers, one has a topological extension of Tverberg theorem [23, 36] which asserts that any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ from the (n1)(n-1)-dimensional simplex to $\mathbb{R}^{d}$ has an $r$-Tverberg tuple – that is, an $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ of non-empty pairwise disjoint faces of $\Delta_{n-1}$ such that $f(\sigma_{1})\cap\cdots\cap f(\sigma_{r})\neq\emptyset$. When $r$ is a prime power, Tverberg’s theorem is recovered by letting $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be a linear map, that is, one which sends each convex combination of vertices of the simplex to the corresponding convex combination of the image of the vertices. In what follows, we identify the vertex set of $\Delta_{n-1}$ with $n]=\{1,\ldots,n\}$ and we let $\operatorname{\mathrm{Vert}}(\sigma)\subseteq[n]$ denote the vertex set of any face $\sigma$ of $\Delta_{n-1}$.
@@
Let $r\geq 2$ be a prime power and let $d,m,n\geq 1$ be integers with $n\geq(r-1)(d+m+1)+1$. Suppose that $\mathcal{F}$ is a family of non-empty subsets of $[n]$ such that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Then for any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$, there exists an $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that no element of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{j})$ for any $j\in[r]$.
As a further example of fan distributions obtainable from Tverberg-type results, we show that the Gale dual of the optimal known results [9] for the “colored Tverberg problem” gives rainbow distributions of a coloring of a finite point set, that is an $r$-fan which distributes the point set so that each of the fan’s open half-flats contains at most one point from each color class.
@@
    Let $r\geq 4$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(d+1)$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{i}|\geq r$ for all $i\in[d+1]$, then there exists a conical $r$-fan in $\mathbb{R}^{n-d-1}$ which rainbow distributes the $(d+1)$-coloring.
@@
    Let $r\geq 2$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(2d+1)$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{2d+1}$ is a $(2d+1)$-coloring of $X$ such that $|X_{i}|\geq r$ for all $i\in[2d+1]$, then there exists a complex regular $r$-fan in $\mathbb{C}^{n-d-1}$ which rainbow distributes the $(2d+1)$-coloring.
As an example of Theorem 1.6, let $r=4$ and $d=1$. Then for any partition of a set of $8$ affinely spanning points in $\mathbb{R}^{6}$ into two color classes of size $4$ each, there is a ($4$-dimensional) conical $4$-fan in $\mathbb{R}^{6}$ whose open half-flats contain at most one point from each color class. If $X$ is typical, then at most one point lies in the fan’s $3$-dimensional center.
@@ 1.4. Distributions by Multiple Fans
Arguably the most fundamental problem in equipartition theory, dating back to Grünbaum [15] and Hadwiger [16], seeks the minimum dimension $d:=\Delta(m,k)$ such that any $m$ finite point sets in $\mathbb{R}^{d}$ can be equipartitioned by the $2^{k}$ regions determined by $k\geq 2$ independent affine hyperplanes. While a degrees of freedom argument  [26] gives the lower bound $\Delta(m,k)\geq\frac{m(2^{k}-1)}{k}$, very few tight upper bounds are currently known. For two hyperplanes in particular, sharp results exist only when either m1m-1, $m$, or $m+1$ is a power of two [8, 7, 20], in which case $\Delta(m,2)=\lfloor 3m/2\rfloor$. When $r$ is an odd prime, we give analogous extensions of our equidistribution results for two fans, where now given an $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a finite point set $X$ in $\mathbb{R}^{d}$ we ask for two $r$-fans $F_{r}^{1}=\cup_{i=1}^{r}B^{1}_{i}$ and $F_{r}^{2}=\cup_{j=1}^{r}B^{2}_{j}$ in $\mathbb{R}^{d}$ whose intersection contains $X$ so that each intersection of the form $\operatorname{\mathrm{Int}}(B_{i}^{1})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})$ contains at most $1/r^{2}$ of the points from each $X_{k}$. As with the calculations of the tight values of $\Delta(m,2)$, our results are ultimately derived from cohomological techniques which impose restrictions on the values of $m$.
@@
Let $d\geq 1$ be an integer, let $r\geq 3$ be a prime number, and let $m\geq 1$ be an integer such that each of the coefficients in the $r$-ary expansion of m(r1)/2m(r-1)/2 is even.
@@
    If $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$, then any $m$-coloring of a set $X$ of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by two conical $r$-fans.
@@
It is straightforward to show (see Remark 5.2) that the $r$-ary expansion of m(r1)/2m(r-1)/2 has all even coefficients provided $m$ is of the form $m=2a(r^{\ell_{1}}+\cdots+r^{\ell_{k}})$ where the $\ell_{i}\geq 0$ are distinct and $1\leq a\leq r-1$ is odd. In particular, one may let $m=2a(r^{\ell}-1)/(r-1)$ for any $\ell\geq 1$. When $r\equiv 1~(\text{mod}~4)$, one may also let $m$ be the sum of distinct powers of $r$, and in particular one may let $m=(r^{\ell}-1)/(r-1)$ for any $\ell\geq 1$.
As was the case for a single fan, Theorem 1.7 is the special case of a Dolnikov-type piercing distribution result for two fans (see Theorem 7.1 below) which in turn is the Gale dual of a linear version of a multiple intersection version of Sarkaria’s theorem. In essence, this result guarantees two Tverberg tuples which are maximally distinct (see  [13] for the corresponding extension of Theorem 1.5 when $r=2$).
@@
Let $d,m,r,n\geq 1$ be integers where $r\geq 3$ is a prime number, the $r$-ary coefficients of m(r1)/2m(r-1)/2 are all even, and $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$. Suppose that $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and that $\mathcal{F}$ is a family of non-empty subsets of $[n]$ such that $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$. Then there exists two Tverberg-tuples $(\sigma_{1},\ldots,\sigma_{r})$ and $(\tau_{1},\ldots,\tau_{r})$ for $f$ such that no member of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{i}\cap\tau_{j})$ for any $i,j\in[r]$.
@@
The remainder of our paper is organized as follows. In Section 2 we briefly discuss Gale duality in the real and complex setting, which we use in Section 3 to establish connections between Tverberg partitions and distributions of point sets by linear fans. This connection, in conjunction with a lifting trick, is used in Section 4 to prove our distribution results for a single fan. Sections  5 and  6 are devoted to the proof of Theorem 1.8, which in Section 7 is used to prove our distribution results for two fans. Finally, in Section 8 we examine typical point sets and prove our sharpness results for equidistributions by a single fan.
@@ 2. Real and Complex Gale Duality
In this section we discuss the properties of the real and complex Gale transform which we will need in translating between Tverberg partitions and fan distributions. These are quite standard in the real setting (see, e.g., [21]) and only require the minor adjustment of taking conjugates to hold in the complex case. Nonetheless, for the sake of completeness we shall include their proofs here.
@@
Now let $\mathbb{K}$ denote the field of real or complex numbers and let $a_{1},\ldots,a_{n}$ be a sequence of $n$ points in $\mathbb{K}^{d}$ which ($\mathbb{K}$-) affinely span $\mathbb{K}^{d}$ (thus $n\geq d+1$). The Gale transform of this sequence is obtained as follows. First, let $A$ be the $(d+1)\times n$ matrix with coefficients in $\mathbb{K}$ whose column vectors are the $\widehat{a_{j}}=(a_{j},1)\in\mathbb{K}^{d+1}$. Because the $a_{j}$ affinely span $\mathbb{K}^{d}$, the $\widehat{a_{j}}$ linearly span $\mathbb{K}^{d+1}$. Thus $A$ has maximal rank and so its kernel has dimension nd1n-d-1. Letting $b_{1},\ldots,b_{n-d-1}$ be a basis for $\ker A$, we let $B$ be the $(n-d-1)\times n$ matrix whose $i$-th row is $b_{i}$, and we let $\overline{B}$ denote its conjugate matrix. The Gale transform of $a_{1},\ldots,a_{n}$ is then defined to be the columns $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ of $\overline{B}$. While the Gale transform is defined only up to a choice of basis for $\ker A$, the basic properties we shall need hold independently of such a choice, and moreover is not difficult to show that a different choice of basis corresponds to a change of the $g_{i}$ by a linear isomorphism of $\mathbb{K}^{n-d-1}$.
@@
Now suppose that $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ satisfies (1) and (2). Let $B$ be the $(n-d-1)\times n$ matrix whose columns are the $g_{i}$ and let $b_{1},\ldots,b_{n-d-1}$ be its rows. As the $g_{i}$ linearly span $\mathbb{K}^{n-d-1}$, so do the $\overline{g_{i}}$ and therefore $\overline{B}$ has maximum rank. Thus $\dim\ker\overline{B}=d+1$. Since $\sum_{i}g_{i}=0$, $\sum_{i}\overline{g_{i}}=0$ as well and so one can choose a basis $w_{1},\ldots,w_{d+1}$ for $\ker\overline{B}$ with $w_{d+1}$ the all one vector. Now define $A$ to be the matrix whose rows are the $w_{i}$. As $A$ has full rank its columns span $\mathbb{K}^{d+1}$, and, since the last row of $A$ is the all one vector, the columns of $A$ are of the form $\widehat{a_{1}}=(a_{1},1),\ldots,\widehat{a_{n}}=(a_{n},1)$. Thus $a_{1},\ldots,a_{n}$ affinely span $\mathbb{K}^{d}$. By construction we have $\text{Row}\,A=\ker\overline{B}$, so $\text{Row}\,\overline{A}=\ker B$ and thus $\ker A=(\text{Row}\,\overline{A})^{\perp}=(\ker B)^{\perp}=\text{Row}\,\overline{B}.$ Hence $\overline{b}_{1},\ldots,\overline{b}_{n-d-1}$ is a basis for $\ker A$, and so the columns $g_{1},\ldots,g_{n}$ of $\overline{\overline{B}}=B$ are the Gale transform of $a_{1},\ldots,a_{n}$. ∎
For our purposes, the essential property of the Gale transform is that it gives a correspondence between affine dependencies among the $a_{1},\ldots,a_{n}$ and allocations of $g_{1},\ldots,g_{n}$ with respect to a codimension one linear subspace. More precisely, let $\langle u,w\rangle=\sum_{i=1}^{n}u_{i}\overline{w}_{i}$ denote the standard Hermitian inner product of two vectors $u$ and $w$ in $\mathbb{K}^{n}$. One then has the following.
@@ 3. Tverberg Partitions and Linear Fans
In this section, we describe an explicit connection between Tverberg $r$-tuples and linear $r$-fans, i.e, $r$-fans whose center is a linear subspace. This connection will be used crucially in the proofs of Theorems 1.11.31.4, and 1.6 in Section 4.
@@
3.1. Linear Conical Fans